If nPr = 336 and nCr = 56, determine what n and r are.
Answer by fcabanski(1391)

Look at the formula for each:


nPr=n!/(n-r)! = 336 and nCr=n!/r!(n-r)! = 56


That means 336/r! = 56 and r! = 336/56 = 6 = 3*2 * 1 so r = 3.


n!/(n-3)! = 336 = n*(n-1)*(n-2) (Remember, if the denominator is (n-3)!, it cancels all but the first 3 factors of n!.


Example 8!/(8-3)! = 8*7*6*5*4*3*2*1/5*4*3*2*1= 8*7*6


The easiest way to finish is to try some numbers.


10: 10*9*8 = 720 (too big)


9: 9*8*7 = 504 (too big)


8: 8*7*6 = 336 (That's it)


n=8 and r = 3
Check my profile and contact me for one on one online tutoring. I can help you understand this with one on one online tutoring. It's one thing to follow an example, but even better if you understand the concept.